While standing at the edge of the roof of a building, a man throws a stone upward with an initial speed of 6.11 ms. The stone subucqoently flls to the ground, which is 149 m below the point where the stonc leaves his hand Ar what spced does the stone impact the ground? Ignore air resistance and use g 9.81 mu/s for the acceleration due to gravity. m/s impact speed How much time is the sione in the air? clapsod time SONY
Answer #1

a)the speed with which the ball was thrown upwards = 6.11 m/s

v^2 - u^2 = 2gH

0 - 6.11^2 = 2*-9.8*H

H = 1.9 m

height to which the ball goes up = 14.9+1.9 = 16.8 m

speed at which the impact = sqrt(2gH) = sqrt(2*9.8*16.8) = 18.14 m/s

b) time taken to reach the top

v = u + gt

0 = 6.11 - 9.8*t

t = 0.623 s

v = u + gt

18.14 = 0 +9.8*t

t = 1.877 s

total time in air = 1.877 + 0.623 = 2.5 s